7 April 2010 UT Technical Night 1

David Gonzalez Huerta, Dave Thompson, John Morris (LBTO) Marco Pedani, Olga Kuhn, Andrew Rakich (Tucson)

Start with LBC

02:00 opening dome; winds are relatively high, about 15 m/s.

02:10 Slewing to flat field position via OB, 2500 counts in 1 sec at Us and Y, scaling by 10 for OB. Decent flats, but on the Red camera we see a ~1000 row long strip at the right of Chip 1. What is it? Looks like something casting a shadow on the chip. This was through Y-FAN. We'll try another filter, but it may be closer.

02:17 Fast OB to take 1 sec tests at V-BESSEL(021914) and r-SLOAN (021905). Running a generic skyflat OB (dec = -90) to see if it dithers (I suspect no but we need to check - I do not see any stars in these flats).

02:33 What is situation with PSF - OK now.

02:40 We are not configured for LBC. (This could explain the strip on the red side chip 1 [_later, no it did not as we saw this band also in Y band data_] and the odd pattern in blue Us flats.)

03:13 Collimating near a pointing star (WT10_271) R(031416) B(031418). Blue needed manual help - we entered -1500 Z7. Then dofpia picked up the pupils, but Z11 and Z22 corrections were not so good.

03:38 Rotator center check: elevation = 69.7 WT10_271
PA = 0 B lbcb.20100407.033700.fits R lbcr.20100407.033657.fits
PA = 180 B lbcb.20100407.033952.fits R lbcr.20100407.033948.fits

PA = 0 B lbcb.20100407.034847.fits R lbcr.20100407.034842.fits
PA = 180 B lbcb.20100407.035059.fits R lbcr.20100407.035056.fits

These datasets give, for the rotator centers of Blue and Red:
Blue: 1032, 2923 (close to what Marco & Dave measured in January, should be updated in headers)
Red: 1080, 2914 (22 pixels off !!! from what it was --- 1080, 2892) *NEW LBC-RED ROTATOR CENTER*

We are going to a different star to check again. WT10_305

04:10 First check pointing --- we are at elevation 84 deg
PA=0: B 041038 R 041035 ptadjust > IE -27.8, CA=-10.7 old values new values IE= -34.6, CA= -38.0 new values (referenced to recently measured Red rotator center 1080, 2914)

04:18 slewing to low elevation star WT10_131, elevation = 34 deg. Collimating there. Initial pupils showed lots of coma but later are OK given the seeing, and after huge focus excursions. However, PSF Right RY appears stuck at RY=8.4". Reported seeing was 1.8 and 2".

Observing pointing star (WT10_131): B 042953 R 042951 run ptadjust to adjust IA... IA = -267.31 old value IA = -359.7 (-92") but...
IE also changed a lot. to -61 from -35.

We are going to make a compromise in IE = -48"

04:35 Another image of WT10_131 with these newly adjusted IA and IE values.

B 043528 R 043522

04:44 Going to ACT0239. Collimated - took about 6 iterations.
PA = 0: B 050327 R 050324
PA = 180: B 050651 R 050648

R again consistent with our earlier measurements tonight: 1078, 2911 for Red Rotator Center

05:12 ACT0307 - again at high elevation. 87.4 deg elevation (incidentally co-pointing looks pretty good)

PA=0 B 052347 R 052344

05:26 Go to LQG field to focus/collimate and observe. Elev = 62 deg. Seeing is estimated to be 1.5" from FPIA. We are observing LQG pointing #1 to check offsets, guiding... Blue is slave, Red master in this OB. Guiding at 8sec R and 4sec Blue. Stellar FWHM 2.1-2.2 " estimated from tech chips.

B 053653 R 053641

053641 - the strip is still there on the right side of chip 1. 054125 - FWHM for Blue ~ 9.3 pxl = 2.1 " 054454 - FWHM for Red ~ 7.8 pxl = 1.7"

06:00 closed because of high winds

08:22 reopening as winds have dropped a little bit. Dave noted that it seemed like a different airmass came in --- temperature had dropped from -1 to -4 early on, but then had risen again and now is about -1 or -2. While closed we decided to spend most of the remainder of the night on LUCIFER - primarily to check the SX collimation which was a problem. But now with this higher temperature, first we'll check IA on an LBC pointing.

Final IA had ended at -382.21.

08:30 ACT0311 at 86deg elevation. Active Optics had been cleared. After this we had hit limits. T~-3 C. David had to fix an actuator. "dofpia" crashed with an error --- it has done this already about 2 times tonight. We need to figure out what this means. Blue side is comatic.
Reading "lbciaCoeffs_red.dat" file ...
 # Error string

Sending data to TCS ...

Checking for Red Convergence
Limits:    400.0   400.0   400.0   400.0   400.0   400.0   100.0
Processing: /newdata/lbcb.20100407.084651.fits
Searching for pupils... 
Computing aberrations using 2 pupils
Focus Offset = -0.80000001  Filter = V-BESSEL
BlueZ22Zero:   2.300
RedZ22Zero:   1.650
Not enough points for estimating Z22
Rejected pupil 0 with diameter 2697.57
Not enough points for estimating Z22
Rejected pupil 1 with diameter 2666.29
% Attempt to subscript RES with <INT      (       3)> is out of range.
% Execution halted at: LBCFPIA           291
  /home/lbcobs/LBCFPIA/lbcfpia/src/lbcfpia.pro
%                      DOFPIA            435
  /home/lbcobs/LBCFPIA/lbcfpia/src/dofpia.pro
%                      LBCFPIA           291
  /home/lbcobs/LBCFPIA/lbcfpia/src/lbcfpia.pro
%                      DOFPIA            387
  /home/lbcobs/LBCFPIA/lbcfpia/src/dofpia.pro
%                      LBCFPIA           291
  /home/lbcobs/LBCFPIA/lbcfpia/src/lbcfpia.pro
%                      DOFPIA            435
  /home/lbcobs/LBCFPIA/lbcfpia/src/dofpia.pro
%                      LBCFPIA           291
  /home/lbcobs/LBCFPIA/lbcfpia/src/lbcfpia.pro
%                      DOFPIA            435
  /home/lbcobs/LBCFPIA/lbcfpia/src/dofpia.pro
%                      $MAIN$          
% Program caused arithmetic error: Floating divide by 0

We have coma - need to give -1500nm Z7 to SX (this is becoming a pattern).

09:09 ACT0311 B 090909 R 090905 new values for CA and IE: CA=-29.6 IE=-39.7 at +80 deg elevation, T=-2.4C

09:13 ACT0105 RY on DX may be stuck again - David cleared it. Note that the earlier instance was also at low elevation (RY sticky).
B 092029 R 092025

IE changed by 30"
IE now -64.5 and CA is almost unchanged, -33. IA almost unchanged -385.61. ambient temperature, T=-2C.

09:30 We are switching now to LUCIFER.

-- OlgaKuhn - 07 Apr 2010
Topic revision: r3 - 07 Apr 2010, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback